# ChemReasoner: Heuristic Search over a Large Language Model’s Knowledge Space using Quantum-Chemical Feedback

### v3

@@ 6 Conclusion
We introduce a multi-modal framework unifying linguistic reasoning enabled by generative LLMs with atomistic structure based rewards, built on principles in catalysis and quantum chemistry. We demonstrate that such integration can enable the recommendations of novel catalyst structures that are critical for developing energy efficient chemical conversion processes and combat climate change. Our methodology is generic and can be applied to other scientific applications in biology and chemistry. Carrying strong quantitative and qualitative validation with domain guided approaches, and extending support towards recommendation of novel structures that are not covered in the original training datasets remain major candidates for future work.
@@ Appendix A GNN Driven Adsorption Energy Calculation
Attempting to capture all possible disambiguation of catlyst is intractable for large tree searches, especially with less-well behaved models such as GPT-3.5-turbo. It also becomes more difficult to decide which catalysts to penalize as containing non-metals, such as Copper/Zinc-oxide. Therefore, we prompt the LLM to parse out the correct chemical symbols with the LLM into digestible lists of chemical symbols. For instance, the above phrases would be parsed into:
@@ C.1 Parameter selection
Given the compute intensive nature of each ChemReasoner query, we carried out hyperparameter sweeps to determine the optimal number of actions for search expansion (), as well as the maximum number of relaxation steps (). Figure 8 shows how increasing impacts the distribution of rewards, and we set it to a default value of 8 for all experiments. Figure 6 shows convergence of the GNN for finding the optimal atomistic structure configuratio over 300 iterations for 40 random initial adsorbate+catalsyst configurations. Relaxations are terminated early if ev/Å. At each step in the relaxation process, the difference between the current adsorption energy and the final adsorption energy is reported. The dashed gray line depicts the iterations cutoff for reward calculations in ChemReasoner, 64 steps. Most iterations are nearly converged by 64 steps.
@@ C.3 Analysis of Search Depth
C.4 Distribution of Reward Functions
C.5 Ablation on Actions Taken

### v4

@@ 6 Conclusion
[…]
Our methodology makes use of instruction-following foundational LLMs, GPT-3.5 and GPT-4, avoiding the need for expensive finetuning. Additionally, our application of automatic planning methods removes the need to hard-code domain specific knowledge into the search method. So long as one can define a systematic search space and a relevant domain specific reward function, one could adapt our method to other scientific applications in biology, chemistry, or materials science. While we recognize the importance of accessible, open-source LLMs, our experiments with Llama-2 revealed that such models may not currently possess the instruction following capabilities or scalability required for our heuristic search. Therefore, the development of open-source LLMs for scientific reasoning remains the subject of future work.
Finally, ChemReasoner presents a new way to carry out computational investigations of catalyst materials with DFT surrogate models. Our application of a GNN based reward function grounds the LLM in reality, guiding the search towards catalysts with higher activity. Still, a limitation of ChemReasoner is the reliance on existing reference structures, which may not accurately reflect real catalyst structures, providing a challenge to verification of results with DFT. This motivates the future development of more robust methods to convert textual representations of materials into 3D atomistic structures.
Acknowledgements
This work is supported in part by the PNNL seed Laboratory Directed Research and Development (LDRD) program, Generative AI for Science, Energy and Security Lab-Wide Laboratory Directed Research and Development Investment.
This work is also in part based upon work supported by the Molecule Maker Lab Institute: an AI research institute program supported by NSF under award No. 2019897. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
PNNL acknowledges the Microsoft Azure Quantum Elements team for their assistance with this study and their contributions to this publication.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning for the chemical and material sciences. It involves the use of a large language model (LLM), which should be treated as a black-box. LLMs are known to “hallucinate”—produce erroneous or fabricated text—and may plagiarise other scientific work. The goal of this paper is to incorporate a domain specific reward function to mitigate hallucinations by the LLM by validating their answers with computational chemistry feedback. Still, scientific results produced by LLMs should still be stringently verified by domain-experts to avoid false claims. The primary focus of this work is to discuss the strengths and weaknesses of the AI system proposed herein as a tool to accelerate scientific discovery, not necessarily to claim the discovery of a novel catalyst.
@@ Appendix A GNN Driven Adsorption Energy Calculation
Attempting to capture all possible disambiguation of catlyst is intractable for large tree searches, especially with less-well behaved models such as GPT-3.5-turbo. It also becomes more difficult to decide which catalysts to penalize as containing non-metals, such as Copper/Zinc-oxide. Therefore, we prompt the LLM to parse out the correct chemical symbols into digestible lists of chemical symbols (see Appendix B.5 for the prompt details). We would expect the above phrases to be parsed into:
@@ B.3 BioFuels Dataset
B.5 Symbols Parsing Prompt
One critical requirement of ChemReasoner’s pipeline is the ability to a fuzzy, linguistic catalyst name into a list of symbols (see Appendix A). To accomplish this task, we use the following LLM prompt template: Consider the following list of catalysts:
[{cat_0}, {cat_1}, ..., {cat_k}].
For each catalyst, return the list of chemical symbols that make up the catalyst. If a catalyst does not have a chemical symbol, return None. If a catalyst is already a chemical formula, repeat the elements in the chemical formula.
Format your list as:
{cat_0}: [list_0]
{cat_1}: [list_1]
{cat_k}: [list_k]
Where cat_i is the ith catalyst proposed by the LLM. The symbols lists can be parsed out from the answer by following the provided format in the prompt.
B.6 Planner Prompt
Here, we detail how ChemReasoner determines the next set of catalyst descriptors by implementing a planner prompt. The prompt incorporates context from previous catalyst answers and the current search state to update the search states in a reasonable way. The search state corresponds to the 4 action types listed in section 5.1:
1.
Catalyst Type: the current catalyst type,
2.
Exclusion Criteria: the list of catalyst design criteria to exclude,
3.
Inclusion Criteria: The set of catalyst criteria to include,
4.
Relationship to previous candidate set: How the new catalyst should relate to the previous catalysts (e.g. similar to, different from).
ChemReasoner uses actions to modify these four state variables. Whereas ChemReasoner-Expert uses a hand-defined set of actions designed by catalysis experts, ChemReasoner-Planner uses the LLM to determine the actions. To this end, we propose the following contextualized prompt template for our planner module, presented in figure 6.
In the planner template, the phrases denoted by $ are treated as “variables” in the LLM’s context window. They hold information about the following information:
1.
search_state: a dictionary mapping each ChemReasoner state variable name to its current value,
2.
action_space: a description of the possible actions,
3.
root_question: the root question for ChemReasoner to answer. For example, “What is a good catalyst for X reaction?”
4.
current_prompt: the previous catalyst prompt that produced the current_answer. It should be a modification of root_question with the previous actions chosen by ChemReasoner,
5.
current_answer: the current catalyst recommendations and their scientific explanations, as returned by the LLM.
@@ C.1 Parameter selection
Given the compute intensive nature of each ChemReasoner query, we carried out hyperparameter sweeps to determine the optimal number of actions (), as well as the maximum number of relaxation steps (). Figure 9 shows how increasing impacts the distribution of rewards, and we set it to a default value of 8 for all experiments. Figure 7 shows convergence of the GNN for finding the optimal atomistic structure configuratio over 300 iterations for 40 random initial adsorbate+catalsyst configurations. Relaxations are terminated early if ev/Å. At each step in the relaxation process, the difference between the current adsorption energy and the final adsorption energy is reported. The dashed gray line depicts the iterations cutoff for reward calculations in ChemReasoner, 64 steps. Most iterations are nearly converged by 64 steps.
@@ C.3 Analysis of Search Depth
C.4 Ablation on Actions Taken